# EEsizer: LLM-Based AI Agent for Sizing of Analog and Mixed Signal Circuit

### v1

@@ I Introduction
The rapid expansion of Machine Learning (ML) techniques has increasingly highlighted its potential for automating circuit design, offering new possibilities for achieving automation in the circuit sizing process. Bayesian Optimization (BO) models transistor sizing as a black-box optimization problem, employing Gaussian Process (GP) regression to efficiently guide the search for optimal solutions [5, 7]. Deep Reinforcement Learning (DRL) further improves sizing efficiency by leveraging rewards from iterative simulations, enabling adaptation to complex design spaces [8]. However, these methods operate without domain-specific analog design knowledge, leading to high computational costs and inconsistent performance across varying circuit designs and performance metrics. [9].
Figure 1: Typical circuit design flow. The process begins with front-end stages including topology selection and circuit sizing, followed by back-end tasks such as placement and routing. Circuit sizing is particularly challenging due to high-dimensional design spaces and complex trade-offs among performance metrics, often requiring iterative simulations and expert intervention.
@@
Method
Sizing
Optimizer
Perf.
Metrics
Node
Transfer
Topology
Gen.
AmpAgent[16]
ABC[17]
TuRBO[18]
4 No Yes
@@
To evaluate the proposed agent, we first benchmarked eight LLMs on six basic circuits, demonstrating the EEsizer’s compatibility across different models, and selected three high-performing models—Gemini 2.0 Flash, GPT-4.1, and OpenAI o3—for further study. We then applied the EEsizer to a 20-transistor CMOS operational amplifier with a complementary input stage and a class-AB output stage. The agent was tasked with achieving ten performance metrics, including rail-to-rail operation, across 180 nm, 130 nm, and 90 nm PTM technology nodes. Finally, we manually designed a bias circuit and performed variation tests on both transistor dimensions and threshold voltages to assess the reliability of the optimized circuit.
@@ II Methodology
The entire process for circuit sizing with EEsizer begins with user input. In the optimization process, action, observation, and comparison create a Reasoning and Acting (ReAct) loop [28], where the agent cycles through reasoning (analyzing) and taking action based on its insights. It starts with sizing prompt generation. The agent then selects appropriate functions based on performance metrics. After executing these functions, it reviews the results and compares them with the user’s input to check for alignment. The final output is produced if the results meet the performance criteria. To improve the agent’s interpretability, this agent will provide not only the modified netlist but also the reasoning behind each parameter adjustment in every iteration. The entire process for circuit sizing with the EEsizer is shown in Fig. 2.
@@ II-E Simulation in-loop
Figure 6: Circuit configurations for measurement: (a) Unity-gain configuration for offset, ICMR. (b) Open-loop configuration for output swing, with the non-inverting input fixed at Vsup/2. (c) Configuration for CMRR. The measurement methods are from [31].
In the function calling process, the agent sends a question to the function, which is the Task 4 result, including netlist and target performance, to the LLM. The LLM selects the appropriate function and arguments, which the agent then uses to execute the necessary simulations via Ngspice, an open-source circuit simulator. After running the simulation, Ngspice returns the results to the agent, which executes additional functions as needed. The agent finally processes the output and converts it into a natural language response, providing the user with a clear and understandable answer. If the user’s question does not require function calls, the model generates a response as usual. In this case, the LLM is called only once.
@@
CMRR Amplifier The minimum cmrr measured across Vcm range Amplifier
@@ II-F Comparison and Iteration
In the convergence check stage, the algorithm determines whether the optimization has successfully met all design targets. Success is declared only if (1) all specification flags are true and (2) no transistors are operating in sub-threshold regions ($v_{gs}-v_{th}<0$), which minimizes leakage currents and improves circuit stability. If success is achieved, the optimization terminates and provides the final results to the user. Otherwise, the feedback, which includes result history and the latest outputs, is used to guide the next iteration of parameter updates. This mechanism ensures a closed-loop process, where simulation results are systematically validated against specifications, preventing invalid operating conditions and promoting reliable convergence toward the final design solution.
@@
if all $s\_pass=\text{True}$ and $vgscheck=$ “No values found where $v_{gs}-v_{th}<0$then
@@ III-A Initial Test
In this work, we extend [27] to evaluate eight LLMs: Anthropic’s Claude 3.5 Sonnet, Claude 3 Sonnet, Claude 3 Haiku; OpenAI’s GPT-4o mini, GPT-4o, GPT-4.1, o3; and Google DeepMind’s Gemini 2.0 Flash across six different circuits. Each model was assessed using identical prompts, functions, and performance metrics. The evaluation focused on the number of iterations required to achieve successful optimization, with a maximum limit of 20 iterations per run. All the tests used a PTM 180nm transistor model and initialized with W=1 µm, L=180 nm. The performance metrics targeted for each circuit are listed in Table. IV and the results are summarized in Fig. 8. OpenAI o3 exhibited the best overall performance, achieving a 100% success rate across all circuits with the fewest iterations and the lowest variability, despite requiring relatively more iterations on the ring oscillator. Claude 3.5 Sonnet also achieved a 100% success rate, however, it required noticeably more iterations on the 5T-OTA circuit. GPT-4.1 performed comparably on most circuits, however, it achieved only a 40% success rate on the ring oscillator. Gemini 2.0 Flash performed worse on the XOR gate, however, it achieved higher success rates on the other circuits, despite exhibiting larger variations. Claude 3 Sonnet and GPT-4o mini exhibited similar trends, with lower success rates and higher variability, particularly for the NAND and XOR gates. Claude 3 Haiku succeeded on digital circuits, however, it consistently failed on the oscillator and 5T-OTA.
@@
Circuit Type
Num.
Transistors
Target Performance
@@
Motivated by these findings, we focus on the best-performing LLMs OpenAI o3 and GPT-4.1, which excelled on the optimization of 5T-OTA, along with the top mid-sized model, Gemini 2.0 Flash, to optimize a complementary input stage and a class-AB output stage consisting of 20 transistors [32]. The schematic is shown in Fig. 9. In all tests, 10 performance metrics, including rail-to-rail operation and load conditions, are required. In the experiment, Gemini 2.0 Flash, GPT-4.1, and OpenAI o3 were separately used to optimize the opamp on PTM 180 nm, 130 nm, and 90 nm technology nodes  [33]. All the tests were performed in August of 2025.
@@ III-C The Performance on Different Technology Nodes
Table V presents the results of opamp optimization across 180 nm, 130 nm, and 90 nm technology nodes by LLM Gemini 2.0 Flash. Each node was tested with five independent optimization attempts starting from the same initial conditions and target specifications. A load capacitor $C_{L}=10\,\text{pF}$ and load resistor $R_{L}=1\,\text{k}\Omega$ were used for all tests, and a 5% tolerance was applied to all performance metrics. Iterations exceeding 25 without meeting the target were considered failures.
@@
TABLE V: Performance evaluation for five attempts across three technology nodes: 180 nm, 130 nm, 90 nm. For all the tests, a load capacitor $C_{L}=10\,\text{pF}$ and load resistor $R_{L}=1\,\text{k}\Omega$ is considered. A 5% tolerance is applied to all metrics. If the 25th iteration is reached without meeting any of the targets, it is marked as a failure.
Node Iter.
Gain
UGBW
PM
Power
CMRR
THD
Offset
Output
Swing
ICMR
@@ III-D The Performance on Different Groups of Metrics
We set 10 performance metrics of operational amplifiers across three configuration groups. In all groups, rail-to-rail operation is required, with targets set for the full input common-mode range and output swing. The groups, however, differ in bandwidth, load conditions, and power consumption requirements to emulate realistic application scenarios. As an illustrative example, tests were conducted using OpenAI o3 on the PTM 90 nm technology. Table VI summarizes the performance metrics for each group, including the final results, number of passed and failed attempts.
TABLE VI: Performance evaluation for five attempts across three groups (G1, G2 and G3). A 5% tolerance is applied to all metrics; deviations beyond this tolerance are marked in red. Different targets for G1 and G2, G1 and G3 are marked in bold. For G1 and G3, a load capacitor $C_{L}=10\,\text{pF}$ and load resistor $R_{L}=1\,\text{k}\Omega$ is considered. For G2, a load capacitor $C_{L}=50\,\text{pF}$ and load resistor $R_{L}=100\,\text{k}\Omega$ is consideredIf the 25th iteration is reached without meeting any of the targets, it is marked as a failure.
Group Iter.
Gain
UGBW
PM
Power
CMRR
THD
Offset
Output
Swing
ICMR
@@
The optimization process of the opamp under G 1 is illustrated in Fig. 11. Although fluctuations are observed across all metrics, the agent continues to refine its strategy and successfully meets the specifications in two out of five attempts within 25 iterations. Notably, input common-mode range, output swing, input offset, and unity-gain bandwidth begin far from the targets, however, they approach the desired range after only a few iterations. This few-shot convergence suggests that the LLM-based agent possesses inherent analog design knowledge, enabling it to adapt transistor sizing effectively with limited iterations. Those already in the target range initially or after a few iterations may still fluctuate or even get worse due to the optimization process prioritizing other performance metrics. This occurs because the agent has not yet achieved a balance between all required performance criteria.
Figure 11: Optimization results for the opamp. (a) Input Common-Mode Range (b) Phase Margin (c) Unity-Gain Bandwidth (d) Output Swing (e) Gain (f) CMRR (g) Input Offset (h) Power (i) THD. Metrics that reach the target early may still fluctuate due to the agent prioritizing other objectives before achieving a global balance across all specifications.
TABLE VII: Transistor size and bias voltages for G1-2, G2-5 and G3-5 after optimization
@@
The transistor sizes and bias voltages belonging to a successful attempt of each configuration group are shown in Table VII.
Figure 12: The complete circuit of Fig. 9, which includes a manually designed bias circuit targeting at providing bias voltages in Table. VII.
Figure 13: Variation results for G1-2 based on 10 samples. A standard Gaussian distribution with $\mu=0$, $\sigma=5$ nm and $\mu=0$, $\sigma=10$ mV, was performed on transistor size and $V_{th}$ respectively. (a) Input error across Vin,cm (b) Spectrum of Transient Output (c), (e) Small-signal Bode plot (d) Transient output (f) CMRR across Vcm.
@@ III-E Variation Test
To further assess the robustness of the optimized design, a bias circuit was manually implemented to provide the bias voltages listed in Table VII. A variation test was then performed on circuit G1-2 from Table VI. In this test, device-level variations were introduced to both transistor dimensions and threshold voltage ($V_{th}$). Transistor sizes were varied according to a Gaussian distribution with $\mu=0$, $\sigma=5$ nm, while $V_{th}$ variations were modelled by adding a voltage source at the gate with $\mu=0$, $\sigma=10$ mV. The complete circuit configuration, including bias circuits, is shown in Fig. 12.
The small-signal and large-signal simulation results under ten variation samples are summarized in Fig. 13. Despite device-level perturbations, DC and transient behaviors such as input offset and output swing remain stable, demonstrating rail-to-rail performance. The largest sensitivity is observed in gain, primarily due to variations in transconductance ($g_{m}$) and parasitic capacitances, which in turn affect the unity-gain bandwidth (UGBW) and overall frequency response.
@@
Figure 14: Variation results for G1-2 based on 50 samples. Variations were applied to transistor size and threshold voltage ($V_{th}$), following standard Gaussian distributions with $\mu=0$, $\sigma=5$ nm and $\mu=0$, $\sigma=10$ mV, respectively. The histograms represent the probability density, while the dashed lines indicate the Gaussian fit for each metric based on the calculated $\mu$ and $\sigma$. The performance that hits the target is considered a pass. The target performance group G1 is listed in Table. VI.
@@ IV Discussion
Second, although a simplified variation test was performed on transistor size and $V_{th}$, a full Monte Carlo simulation with accurate device models incorporating Monte Carlo parameters should also be used to rigorously evaluate design robustness. As shown in Fig. 14, not all performance metrics achieved a 100% pass rate, highlighting the need to integrate Monte Carlo simulation into the agent workflow as well.
@@
Finally, optimization in our experiments was constrained to 25 iterations and five trials per configuration because of API rate and cost limits. These practical limits restrict the empirical coverage and make it difficult to fully characterize model behavior. Importantly, simply increasing the iteration budget does not guarantee better results: the LLM can oscillate between candidate solutions or repeatedly explore similar regions of the design space. Large, reasoning-focused LLMs such as OpenAI o3 tend to produce higher-quality strategies and more stable proposals, however, their inference cost and latency make them unsuitable for dense, closed-loop experimentation over many trials.

### v2

@@ I Introduction
The rapid expansion of Machine Learning (ML) techniques has increasingly highlighted its potential for automating circuit design, offering new possibilities for achieving automation in the circuit sizing process. Bayesian Optimization (BO) frames transistor sizing as an optimization problem in which the objective function is evaluated through simulation, and uses Gaussian Process (GP) regression to efficiently guide the search toward high-performing solutions [5, 7]. Deep Reinforcement Learning (DRL) further improves sizing efficiency by leveraging rewards from iterative simulations, enabling adaptation to complex design spaces [8]. However, these methods operate without domain-specific analog design knowledge, leading to high computational costs and inconsistent performance across varying circuit designs and performance metrics. [9].
Figure 1: A typical circuit design flow. The process begins with front-end stages including topology selection and circuit sizing, followed by back-end tasks such as placement, routing, and post-layout simulation. The results from post-layout simulation often necessitate adjustments to the circuit sizing.
@@
Method Sizing Optimizer Perf. Metrics Node Transfer Topology Gen.
AmpAgent[16] ABC[17] TuRBO[18] 4 No Yes
@@
To evaluate the proposed agent, we first benchmarked eight LLMs on six basic circuits, demonstrating the EEsizer’s compatibility across different models, and selected three high-performing models—Gemini 2.0 Flash, GPT-4.1, and OpenAI o3—for further study. We then applied the EEsizer to a 20-transistor CMOS operational amplifier with a complementary input stage and a class-AB output stage. The agent was tasked with achieving ten performance metrics, including rail-to-rail operation, across 180 nm, 130 nm, and 90 nm on Predictive Technology Model (PTM) technology nodes [28]. Finally, we manually designed a bias circuit and performed variation tests on both transistor dimensions and threshold voltages to assess the reliability of the optimized circuit.
@@ II Methodology
The entire process for circuit sizing with EEsizer begins with user input. In the optimization process, action, observation, and comparison create a Reasoning and Acting (ReAct) loop [29], where the agent cycles through reasoning (analyzing) and taking action based on its insights. It starts with sizing prompt generation. The agent then selects appropriate functions to call external simulator and analyze simulation result based on performance metrics. After executing these functions, it reviews the results and compares them with the user’s input to check for alignment. The final output is produced if the results meet the performance criteria. Ngspice, an open-source SPICE simulator, is integrated into EEsizer to enable in-loop simulation, ensuring transparency and re-producibility through the public release of our source code and benchmark data. The transistor models used in this work are PTM models [28], which are open-source models used for academic and circuit design research. To improve the agent’s interpretability, this agent will provide not only the modified netlist but also the reasoning behind each parameter adjustment in every iteration. The entire process for circuit sizing with the EEsizer is shown in Fig. 2.
@@ II-E Simulation in-loop
Figure 6: Circuit configurations for measurement: (a) Unity-gain configuration for offset, ICMR. (b) Open-loop configuration for output swing, with the non-inverting input fixed at Vsup/2. (c) Configuration for CMRR measurement, sweeping Vcm from 0 to VDD.
In the function calling process, the agent sends a question to the function, which is the Task 4 result, including netlist and target performance, to the LLM. After running the simulation, Ngspice returns the results to the agent, which executes additional functions as needed. The agent finally processes the output and converts it into a natural language response, providing the user with a clear and understandable answer. If the user’s question does not require function calls, the model generates a response as usual. In this case, the LLM is called only once.
@@
CMRR Amplifier The minimum CMRR measured across the whole Vcm range Amplifier
@@ II-F Comparison and Iteration
if all $s\_pass=\text{True}$ and $vgscheck=$ “No values found where $V_{\text{gs}}-V_{\text{th}}<0$then
@@
In the convergence check stage, the algorithm determines whether the optimization has successfully met all design targets. Success is declared only if (1) all specification flags are true and (2) no transistors are operating in sub-threshold regions ($V_{\text{gs}}-V_{\text{th}}<0$), which is constrained to strong inversion reduce the effect of variations and the need for extensive Monte Carlo simulations within the optimization loop. If success is achieved, the optimization terminates and provides the final results to the user. Otherwise, the feedback, which includes result history and the latest outputs, is used to guide the next iteration of parameter updates. This creates a dynamic priority system, where the feedback for each iteration is inherently focused on the specifications that have most recently failed, guiding the LLM directly toward the remaining design gaps, thereby converging to a final design solution.
@@ III-A Initial Test
In this work, we extend [27] to evaluate eight LLMs: Anthropic’s Claude 3.5 Sonnet, Claude 3 Sonnet, Claude 3 Haiku; OpenAI’s GPT-4o mini, GPT-4o, GPT-4.1, o3; and Google DeepMind’s Gemini 2.0 Flash across six different circuits. Each model was assessed using identical prompts, functions, and performance metrics. The evaluation focused on the number of iterations required to achieve successful optimization, with a maximum limit of 20 iterations per run. All the tests used a PTM 180nm transistor model and initialized with W=1 µm, L=180 nm. All the tests were performed in August of 2025.
The performance metrics targeted for each circuit are listed in Table. IV and the results are summarized in Fig. 8. OpenAI o3 exhibited the best overall performance, achieving a 100% success rate across all circuits with the fewest iterations and the lowest variability, despite requiring relatively more iterations on the ring oscillator. Claude 3.5 Sonnet also achieved a 100% success rate, however, it required noticeably more iterations on the 5T-OTA circuit. GPT-4.1 performed comparably on most circuits, however, it achieved only a 40% success rate on the ring oscillator. Gemini 2.0 Flash performed worse on the XOR gate, however, it achieved higher success rates on the other circuits, despite exhibiting larger variations. Claude 3 Sonnet and GPT-4o mini exhibited similar trends, with lower success rates and higher variability, particularly for the NAND and XOR gates. Claude 3 Haiku succeeded on digital circuits, however, it consistently failed on the oscillator and 5T-OTA.
@@
Circuit Type Num. Transistors Target Performance
@@ III-B The Performance of Different LLMs on the OpAmp
Motivated by the above findings, we focus on the best-performing LLMs OpenAI o3 and GPT-4.1, which excelled on the optimization of 5T-OTA, along with the top mid-sized model, Gemini 2.0 Flash, to optimize a complementary input stage and a class-AB output stage consisting of 20 transistors  [32]. The schematic is shown in Fig. 9. In all tests, 10 performance metrics, including rail-to-rail operation and load conditions, are required. In the experiment, Gemini 2.0 Flash, GPT-4.1, and OpenAI o3 were separately used to optimize the opamp on PTM 180 nm, 130 nm, and 90 nm technology nodes. All the tests were performed in August of 2025.
@@
TABLE V: Computation efficiency Comparison between Different LLMs.
Model Average Iterations Total Optimization Time Optimization Time per Iteration Simulation Time per Iteration
OpenAI o3 22 73m 27.43s 3m 15.83s 3.98s
GPT 4.1 25 60m 35.26s 2m 27.81s 4.01s
Gemini 2 25 38m 44.08s 1m 32.96s 3.90s
@@
As shown in Table V, the computation time is dominated by the LLM reasoning and communication overhead during each optimization loop. The simulation time per iteration remains nearly constant across all models, indicating that SPICE-level simulation contributes minimally to the overall computation time. Among the evaluated models, Gemini 2.0 achieves the shortest total optimization time, demonstrating high computational efficiency but a lower success rate. In contrast, OpenAI o3 requires the longest total time yet delivers the most reliable optimization results. GPT-4.1 lies between these two, maintaining a balance between reasoning thoroughness and speed. It is important to note that these performance characteristics are dependent on the specific LLM provider and are subject to change with future model updates.
@@ III-C The Performance on Different Technology Nodes
Table VI presents the results of opamp optimization across 180 nm, 130 nm, and 90 nm technology nodes by LLM Gemini 2.0 Flash. Each node was tested with five independent optimization attempts starting from the same initial conditions and target specifications. A load capacitor $C_{\text{L}}=10\,\text{pF}$ and load resistor $R_{\text{L}}=1\,\text{k}\Omega$ were used for all tests, and a 5% tolerance was applied to all performance metrics. Iterations exceeding 25 without meeting the target were considered failures.
@@
TABLE VI: Performance evaluation for five attempts across three technology nodes: 180 nm, 130 nm, 90 nm. For all the tests, a load capacitor $C_{L}=10\,\text{pF}$ and load resistor $R_{L}=1\,\text{k}\Omega$ is considered. A 5% tolerance is applied to all metrics. If the 25th iteration is reached without meeting any of the targets, it is marked as a failure.
Node Iter. Gain (dB) UGBW (MHz) PM () Power (mW) CMRR (dB) THD (dB) Offset (mV) Output Swing (V) ICMR (V)
@@ III-D The Performance on Different Groups of Metrics
We set 10 performance metrics of operational amplifiers across three configuration groups. In all groups, rail-to-rail operation is required, with targets set for the full input common-mode range and output swing. The groups, however, differ in bandwidth, load conditions, and power consumption requirements to emulate realistic application scenarios. As an illustrative example, tests were conducted using OpenAI o3 on the PTM 90 nm technology. Table VII summarizes the performance metrics for each group, including the final results, number of passed and failed attempts.
TABLE VII: Performance evaluation for five attempts across three groups (G1, G2 and G3). A 5% tolerance is applied to all metrics; deviations beyond this tolerance are marked in red. Different targets for G1 and G2, G1 and G3 are marked in bold. For G1 and G3, a load capacitor $C_{L}=10\,\text{pF}$ and load resistor $R_{L}=1\,\text{k}\Omega$ is considered. For G2, a load capacitor $C_{L}=50\,\text{pF}$ and load resistor $R_{L}=100\,\text{k}\Omega$ is consideredIf the 25th iteration is reached without meeting any of the targets, it is marked as a failure.
Group Iter. Gain (dB) UGBW (MHz) PM () Power (mW) CMRR (dB) THD (dB) Offset (mV) Output Swing (V) ICMR (V)
@@
The optimization process of the opamp under G1 is illustrated in Fig. 11. Although performance metrics fluctuate, the agent successfully refined its strategy to meet all specifications in two out of five attempts. The variation across trials arises from the dynamic interaction between the LLM’s probabilistic reasoning and the circuit simulation feedback. In each optimization round, the outcome of the performance calculations guides the LLM’s subsequent parameter adjustments. Failed trials typically occur when the agent explores suboptimal regions of the design space or makes fewer effective decisions during exploration.
Figure 11: Optimization results for the opamp. (a) Input Common-Mode Range (b) Phase Margin (c) Unity-Gain Bandwidth (d) Output Swing (e) Gain (f) CMRR (g) Input Offset (h) Power (i) Total Harmonic Distortion.
TABLE VIII: Transistor size and bias voltages for G1-2, G2-5 and G3-5 after optimization
@@
The transistor sizes and bias voltages belonging to a successful attempt of each configuration group are shown in Table VIII.
Figure 12: The complete circuit of Fig. 9, which includes a manually designed bias circuit targeting at providing bias voltages in Table. VIII.
Figure 13: Variation results for G1-2 based on 10 samples. A standard Gaussian distribution with $\mu=0$, $\sigma=5$ nm and $\mu=0$, $\sigma=10$ mV, was performed on transistor size and $V_{\text{th}}$ respectively. (a) Input error across Vin,cm (b) Spectrum of Transient Output (c), (e) Small-signal Bode plot (d) Transient output (f) CMRR across Vcm.
@@ III-E Variation Test
To further assess the robustness of the optimized design, a bias circuit was manually implemented to provide the bias voltages listed in Table VIII. A variation test was then performed on circuit G1-2 from Table VII. In this test, device-level variations were introduced to both transistor dimensions and threshold voltage ($V_{\text{th}}$). Transistor sizes were varied according to a Gaussian distribution with $\mu=0$, $\sigma=5$ nm, while $V_{\text{th}}$ variations were modelled by adding a voltage source at the gate with $\mu=0$, $\sigma=10$ mV. The complete circuit configuration, including bias circuits, is shown in Fig. 12.
The small-signal and large-signal simulation results under ten variation samples are summarized in Fig. 13. Despite device-level perturbations, DC and transient behaviors such as input offset and output swing remain stable, demonstrating rail-to-rail performance. The largest sensitivity is observed in gain, primarily due to variations in transconductance ($g_{\text{m}}$) and parasitic capacitances, which in turn affect the unity-gain bandwidth (UGBW) and overall frequency response.
@@
Figure 14: Variation results for G1-2 based on 50 samples. Variations were applied to transistor size and threshold voltage ($V_{\text{th}}$), following standard Gaussian distributions with $\mu=0$, $\sigma=5$ nm and $\mu=0$, $\sigma=10$ mV, respectively. The histograms represent the probability density, while the dashed lines indicate the Gaussian fit for each metric based on the calculated $\mu$ and $\sigma$. The performance that hits the target is considered a pass. The target performance group G1 is listed in Table. VII.
@@ IV Discussion
Second, although a simplified variation test was performed on transistor size and $V_{\text{th}}$, a full Monte Carlo simulation with accurate device models incorporating Monte Carlo parameters should also be used to rigorously evaluate design robustness. As shown in Fig. 14, not all performance metrics achieved a 100% pass rate, highlighting the need to integrate Monte Carlo simulation into the agent’s workflow. This integration, will also eliminate the strong inversion requirement used in optimization loop to decide the current results pass or not in Section II-F. Furthermore, future studies should include noise performance. The PTM model we used may not include noise parameters, making accurate noise simulation infeasible. An accurate noise model will enable for an input referred noise and SNR performance measurement.
@@
Finally, optimization in our experiments was constrained to 25 iterations and five trials per configuration because of API rate and cost limits. These practical limits restrict the empirical coverage and make it difficult to fully characterize model behavior. Importantly, simply increasing the iteration budget does not guarantee better results: the LLM can oscillate between candidate solutions or repeatedly explore similar regions of the design space. Large, reasoning-focused LLMs such as OpenAI o3 tend to produce higher-quality strategies and more stable proposals, however, their inference cost and latency make them unsuitable for dense, closed-loop experimentation over many trials. Furthermore, the success rate for the 20-transistor opamp appears low, due to the inherent complexity of the design constraints, the circuit topology, and LLM’s limited expertise in analog circuit design. Although the success rate appears low, in the end, only one successful iteration is needed to achieve the goal of the sizing optimization.